6. For the quadratic function y = -2x^2 + 8x + 1:
a. What are the coordinates (values of x and y) at the vertex of the parabola
b. Does this parabola open upward or downward? Explain.

Answer by fractalier(2101)

Given the quadratic function y = -2x^2 + 8x + 1, we know that the x-coordinate of the vertex lies at -b/2a, or
-8/[2(-2)] = 2
The y-coordinate can be found via f(2)...
-2(2^2) + 8(2) + 1 =
-8 + 16 + 1 = 9
Thus the vertex is at (2, 9).
Since the coefficient of the x^2 term is negative, the graph is concave downward.